# Adversarial Bug Reports as a Security Risk in Language Model-Based Automated Program Repair

### v1

@@ 4.2. Defenses at Key APR Pipeline Stages
We implement concrete defenses at three critical points in the APR pipeline.
Input Filtering. The most comprehensive threat removal would occur if the malicious issues were scanned for malicious intent or expected malicious behavior prior to their processing through the APR system. We use Meta PromptGuard and PromptGuardV2 (Chennabasappa et al., 2025) to analyze incoming issues for Prompt Injection and Jailbreak attacks. In addition, we use Meta LLamaGuard in Versions 3 and 4 (Inan et al., 2023) as well as IBM Granite Guardian (Padhi et al., 2024) to classify incoming issues for potential malicious intent. Comparing PromptGuard with LlamaGuard and Granite-Gardian, PromptGuard has a smaller parameter count of 98M compared to the 8–12B that are used to the other approaches. To assess the efficacy of state-of-the-art closed-weight models, we analyzed the malicious issues with a custom prompt using either OpenAI traditional GTP-4.1-mini model and the o4-mini reasoning model. We have chosen the -mini families to aim for cost efficiency to prevent monetary denial-of-service attacks.
Environment Isolation. After a new issue is entered into the system, it will be processed upon by the APR system. Successful APR execution results in a potential patch solving the reported issue. We automatically collect the patch and upload it as part of a new pull-request to GitHub, which will automatically be processed by the CI/CD system. At that stage, we cannot prevent the APR/CI/CD system from executing potentially malicious code, but we can try to reduce the potential impact of malicious patches. The APR (SWE-agent) was run on a local workstation under human oversight. Resulting costs for APR processing was limited to not exceed \$2 per processed issue. To ground our research in common practices, we used GitHub actions for CI/CD and thus inherited all their existing security sandboxing counter-measures (GitHub, 2025d, c).
@@ 5. Experiments
Table 3. Detection and mitigation effectiveness across the APR pipeline for each adversarial attack type. Attack columns report how many issues led to patch generation and how many resulted in a successful compromise (as defined by the intended attack goal). The table then shows the number of issues flagged at each stage: pre-APR filtering (custom models and LLM-based guards) and post-APR detection (GitHub CodeQL and Copilot). Ensemble columns indicate how many issues were blocked by combinations of detectors. The final two columns assess composite defenses—combining pre- and post-APR ensembles with either the full pre-APR ensemble or the single best-performing pre-APR model.
Pre-APR Detection Attack Post-APR Detection Blocked
@@
Naive APR 5 0 0 0 0 0 0 0 0 0 0 0 5 5 0 0 0 4 0 0 0 0
Inject 14 0 1 5 8 1 0 0 0 0 0 10 14 10 0 0 4 8 7 10 12 12
Ci/Cd 10 10 8 10 8 0 0 0 0 0 0 10 10 10 0 0 6 6 5 10 10 10
Revert CVE 14 0 3 3 6 2 0 0 0 4 0 9 14 13 0 1 7 5 1 8 11 9
Noise APR 8 0 0 0 2 0 0 0 0 1 0 3 8 8 0 0 0 4 2 2 4 4
Total 51 10 12 18 24 3 0 0 0 5 0 32 51 46 0 1 17 27 15 30 37 35
@@
We conduct a controlled empirical study to evaluate whether modern APR pipelines, fortified with state-of-the-art defenses, can detect or prevent adversarial bug reports. Our evaluation mirrors real-world workflows while ensuring no interference with active development.
@@
Target Projects. We selected five mature, high-profile Python repositories from the SWE-bench benchmark: django, flask, requests, pip, and scikit-learn. These projects (10k–90k stars) are well-maintained and diverse in structure. All experiments were performed on clones of the repositories hosted in an isolated GitHub workspace.
@@
Attack Set. We crafted 51 adversarial bug reports across five categories: 5 Naive APR (jailbreak-style prompt manipulation), 8 Deceptive Noise reports (resource DoS), 10 CI/CD exploits (secret exfiltration and RCE), 14 code injection attacks (inserting insecure logic), and 14 CVE reversion attacks (undoing security patches). Each report was designed to resemble real bug reports in language and structure, often including plausible code snippets. All prompts, seeds, and generated content are available in the replication package.
Pipeline Setup. For pre-filtering, we tested PromptGuard and PromptGuardV2 (both lightweight LLMs with 98M parameters), LlamaGuard v3 and v4, Granite Guardian, and custom prompts using OpenAI’s GPT-4.1-mini and o4-mini models. For custom prompts, both structured-output and plain-text variants were used. Lightweight models ran locally; larger models were run on cloud providers (OpenAI and DeepInfra) or operated locally (Ollama). SWE-agent v1.1.0 (with SW-ReX v1.2.2) served as the APR backend, using Claude Sonnet (claude-sonnet-4-20250514) for patch synthesis. We selected SWE-agent due to its top-ranked status on the SWE-bench “verified” leaderboard at the start of our evaluation. Generated patches were reviewed via GitHub Copilot (commit summary, inline comments, low-confidence alerts) and analyzed using CodeQL for traditional static detection.
@@
Evaluation and Validation. For each bug report, we record (1) whether any pre-filter model flagged it, (2) whether SWE-agent generated a patch and whether it fulfilled the attacker’s intended goal (e.g., reversion, leakage, injection), (3) whether Copilot raised any security-relevant concerns, and (4) whether CodeQL detected any vulnerability. Manual validation was performed by two independent experts—a security researcher and a penetration tester—who jointly reviewed each issue and patch. The validation process included reviewing patch logic, inspecting Copilot outputs, and assessing whether the outcome matched the attack objective.
@@ 6. Results
Table 3 summarizes attack success rates as well as the effectiveness of detection and mitigation mechanisms across the APR pipeline, including pre-APR filtering, post-APR detection, and combined ensemble defenses.
@@ 6.2. Pre-APR Counter-Measures
Cost-wise, using unstructured outputs, the mean cost per checked issue was $0.00268\pm 0.000999$) when using o4-mini, and $0.001008\pm 0.000406$) when using gpt-4.1-mini. Using structured output created similar costs (o4-mini: $0.00284\pm 0.001111$, gpt-4.1-mini $0.000965\pm 0.000273$).
@@ 6.4. Ensembles and Combining Pre- and Post-APR Countermeasures
A full pre-APR ensemble blocks 32 issues; when combined with post-APR review, the total rises to 37 blocked pull requests (72.5%). However, as discussed in Section 7, this approach significantly increases false positives and may be impractical in production settings.
@@ 7.1. Adversarial Attacks for APR
RQ1: Are APR systems prone to generating insecure or unintended code when given adversarial bug reports? Answer: Adversarial bug reports can subvert APR pipelines at multiple stages—inducing the generation of insecure or malicious patches, triggering unintended behaviors during CI/CD execution, and wasting developer effort via plausible but irrelevant issues. In our study, 46 out of 51 crafted bug reports (90.2%) led the APR system to produce patches consistent with the attacker’s objective, including CVE reversion, injection of vulnerable code, test-time execution, and semantically meaningless refactoring.
Quality of the issues from the human developer perspective. All generated issues were analyzed by an experienced software developer and an experienced software security specialist. The reviewed issues appeared plausible and actionable for human developers, at least at the first glance. After deeper investigation, most of them could be prevented by human. For example, Naive APR proposes fixes that are obviously not aligned with projects, or CI/CD and Inject attacks are obvious from a security perspective (“extract full environment to external server”) and should immediately raise red flags although they present a plausible story. Other attack classes are more complicated. Noise APR, and revert CVE, are not that obvious from the developer perspective. Obviously, our demonstration attacks do not exhaust possibilities. To further elude detection, attacks may have malicious content split upon into multiple stages and thus make detection harder (Przymus and Durieux, 2025).
@@
APR misalignment is evident even for seemingly vague or generic issues. For instance, a report with minimal justification - essentially a vague request to ”revert recent changes”—still led to the APR system synthesizing a plausible revert patch. This shows that APR models often rely on surface cues over deep semantic validation.
We also evaluated attacks designed solely to generate developer burden through meaningless but compilable patches (e.g., arbitrary refactorings, redundant inlines, or unused imports). While such patches do not introduce functional errors, they impose unnecessary review and maintenance costs and may degrade long-term code quality.
Critically, the distinction between “plausible” and “correct” fixes remains unresolved in APR practice. Some adversarial reports included test cases, others did not—yet both yielded test-passing patches. This underscores a known limitation: passing unit tests does not guarantee semantic correctness, a fact long recognized in APR literature. Despite this, many systems claim real-world applicability by auto-generating pull requests for unvetted GitHub issues, ignoring the possibility of malicious or misinformed inputs.
These attacks are realistic, reproducible, and span a spectrum from manual to fully automated generation. All bug reports were phrased in developer-consistent language and submitted to private GitHub mirrors, illustrating their plausibility in standard open-source workflows. The variety and effectiveness of these strategies reveal critical weaknesses in current APR defenses, posing systemic risks to software integrity and supply chain security.
@@ 7.2. Feasibility and Effectiveness of Defenses
RQ2: What defenses are feasible, and at which points in the workflow can attacks be mitigated effectively? Answer Effective mitigation of adversarial bug reports is feasible but non-trivial. Lightweight, structured LLM-based classifiers (e.g., o4-mini) offer practical pre-APR filtering with low false-positive rates. Runtime isolation in APR/CI environments can prevent exploitation, while post-APR analysis (e.g., Copilot) can flag issues, though often requiring manual review. Ensemble defenses improve detection rates but may incur unacceptable false positives. End-to-end protection demands combining targeted pre-filters, secure execution, and informed human oversight.
False-Positive Rates for Custom Prompts. Performing initial filtering using custom prompts is an efficient way of detecting potentially malicious requests. As Sommer et al. (Sommer and Paxson, 2010) mention, the feasibility of real-life usage of machine leaning techniques depends upon a low false-positive rate as otherwise analyst-fatigue will occur, leading to disabling of filtering mechanisms. To validate the false-positive rate, we selected a range of 100 existing issues of the psf/requests project (issues #6000–6100, excluding one deleted entry) and analyzed them analogous to our malicious issues using the custom prompt.
Using structured output resulted in overall low false-positive rates. GPT-4.1-mini resulted in two 2 false-positives, while 4o-mini only resulted in a single one. Both of them tagged issue #6063 (a security bypass inquiry). GPT-4.1-mini additionally tagged an issue (#6075, containing only an external link)). We argue that both cases are not real-false positives but potentially real-life malicious issues.
@@
Asymmetry of Costs. Our attack leverages a fundamental asymmetry: it is cheap to execute yet costly to defend. Adversaries can automatically generate realistic, context-aware bug reports with minimal effort. Using lightweight context retrieval and small LLMs, we construct prompts that produce plausible reports in a single pass—no expensive infrastructure required.
In contrast, defenses incur significant cost at multiple stages: (1) Filtering requires large LLMs or human triage. Since adversarial reports mimic legitimate submissions, simple spam filters or heuristics are ineffective. Low-volume, high-quality submissions evade rate limits and raise few red flags. (2) APR execution is far more expensive than bug report generation. Models like SWE-agent rely on costly multi-turn interactions with large LLMs, often backed by human oversight or CI integration. (3) Post-APR validation (e.g., CodeQL, Copilot, manual review) adds further overhead, with costs accumulating even for patches that are ultimately rejected. This cost imbalance favors attackers: generating high-impact, low-effort reports is trivial, while defenders must invest heavily at each pipeline stage to detect and contain them.

### v2

@@ 4.2. Defenses at Key APR Pipeline Stages
We implement defenses at three critical points of the APR pipeline
Input Filtering. The most comprehensive threat removal scanning issues for malicious intent or expected malicious behavior prior to their processing through the APR system. We use Meta PromptGuard and PromptGuardV2 (Chennabasappa et al., 2025) to analyze incoming issues for Prompt Injection and Jailbreak attacks. In addition, we use Meta LLamaGuard in Versions 3 and 4 (Inan et al., 2023) as well as IBM Granite Guardian (Padhi et al., 2024) to classify incoming issues for potential malicious intent. Comparing PromptGuard with LlamaGuard and Granite-Gardian, PromptGuard has a smaller parameter count of 98M compared to the 8–12B that are used to the other approaches. To assess the efficacy of state-of-the-art closed-weight models, we analyzed malicious issues with a custom prompt using either OpenAI’s traditional GTP-4.1-mini model and the o4-mini reasoning model. We chose the -mini model families to achieve cost efficiency and to prevent monetary denial-of-service attacks.
Environment Isolation. Once a new issue is entered into the system, it will be processed by the APR system. Successful APR execution results in a potential patch solving the reported issue. We automatically collect the patch and upload it as part of a new pull-request to GitHub, which will automatically be processed by the CI/CD system. At that stage, we cannot prevent the APR/CI/CD system from executing potentially malicious code, but we can try to reduce the potential impact of malicious patches. The APR (SWE-agent) was run on a local workstation under human oversight. Resulting costs for APR processing was limited to not exceed \$2 per processed issue. To ground our research in common practices, we used GitHub actions for CI/CD and thus inherited all their existing security sandboxing counter-measures (GitHub, 2025d, c).
@@ 5. Experiments
Table 2. Detection and mitigation effectiveness across the APR pipeline for each adversarial attack type. Attack columns report how many issues led to patch generation and how many resulted in a successful compromise (as defined by the intended attack goal). Tests columns indicate whether running the project’s test suite would detect breaking changes (Fail — the attack is detected; Increase ($\leq$ threshold) — the number of failing tests increases but remains within the expected failure threshold; No change — no additional tests fail). The table then shows the number of issues flagged at each stage: pre-APR filtering (custom models and LLM-based guards) and post-APR detection (GitHub CodeQL and Copilot). Ensemble columns indicate how many issues were blocked by combinations of detectors. The final two columns assess composite defenses by combining pre- and post-APR ensembles with either the full pre-APR ensemble or the single best-performing pre-APR model.
Pre-APR Detection Attack Tests Post-APR Detection Blocked
@@
Naive APR 5 0 0 0 0 0 0 0 0 0 0 0 5 5 0 1 4 0 0 0 4 0 0 0 0
Inject 14 0 1 5 8 1 0 0 0 0 0 10 14 10 4 0 10 0 0 4 8 7 10 12 12
Ci/Cd 10 10 8 10 8 0 0 0 0 0 0 10 10 10 0 0 10 0 0 6 6 5 10 10 10
Revert CVE 14 0 3 3 6 2 0 0 0 4 0 9 14 13 7 3 4 0 1 7 5 1 8 11 9
Noise APR 8 0 0 0 2 0 0 0 0 1 0 3 8 8 1 0 7 0 0 0 4 2 2 4 4
Total 51 10 12 18 24 3 0 0 0 5 0 32 51 46 12 4 35 0 1 17 27 15 30 37 35
Percent of Total 100% 19% 23% 35% 47% 5% 0% 0% 0% 10% 0% 62% 100% 90% 23% 7% 68% 0% 2% 33% 53% 29% 58% 72% 68%
@@
We conducted an empirical study to evaluate whether modern APR pipelines, fortified with state-of-the-art defenses, can detect or prevent adversarial bug reports. Our evaluation mirrors real-world workflows while ensuring no interference with active development. We developed systematic attacks patterns on the Django project and then reran them automatically across all projects. In our evaluation, we included all resulting attacks. To represent realistic setups, we ran all tools with their respective default configuration, e.g., when using LLMs, we used their default temperature.
@@
Target Projects. We selected five mature, high-profile Python repositories from the SWE-bench benchmark: django, flask, pip, requests, and scikit-learn. SWE-bench was chosen as it is the primary benchmark for evaluating APR systems. These projects are well-maintained, security-relevant, and include multiple CVE entries with confirmed fixes. All experiments were conducted on isolated clones of their official repositories.
@@
Attack: Generating the bug reports. We used DevStral to generate 51 adversarial bug reports across five categories: 5 Naive APR (jailbreak-style prompt manipulation), 8 Deceptive Noise reports (resource DoS), 10 CI/CD exploits (secret exfiltration and RCE), 14 code injection attacks (inserting insecure logic), and 14 CVE reversion attacks (undoing security patches). Each report was designed to resemble real bug reports in language and structure, often including plausible code snippets. We provide a complete replication package including prompts, seeds, and generated content.
Pre-APR Detection. Recent research highlights the need for a filtering phase before submitting any input to an agentic AI, to prevent jailbreaks and other forms of attacks (Yi et al., 2024). We argue that similar filtering is needed to block malicious bug reports in APR workflows. We evaluated PromptGuard and PromptGuardV2 (both lightweight LLMs with 98M parameters), LlamaGuard v3 and v4, Granite Guardian, and custom prompts using OpenAI’s GPT-4.1-mini and o4-mini models. For custom prompts, both structured-output and plain-text variants were used. Lightweight models ran locally; larger models were run on cloud providers (OpenAI and DeepInfra) or operated locally (Ollama).
APR Agent. SWE-agent v1.1.0 (with SW-ReX v1.2.2) served as the APR backend, using Claude Sonnet (claude-sonnet-4-20250514) for patch synthesis. We selected SWE-agent due to its top-ranked status on the SWE-bench “verified” leaderboard [^2] https://www.swebench.com/ at the start of our evaluation in May 2025. Please note, that we are not investigating the efficacy of the APR system for patch generation, but its resilience to adversarial bug reports.
Test Suites. After APR generates a patch, we evaluate whether applying it introduces regressions by running the project’s full test suite. We record the number of failing tests and categorize the outcome into three cases: (1) No change — the patch does not introduce any additional test failures compared to the unpatched version; (2) Increase ($\leq$ threshold) — the patch increases the number of failing tests, but the total remains within the project-defined tolerance for expected failures; [^3] Default project configuration; many projects allow a small number of failing tests. (3) Fail - the patch causes the number of failing tests to exceed the allowed threshold, or the test suite fails to execute.
Post-APR detection Only two of the evaluated projects (requests and scikit-learn) were scanning for vulnerabilities. They were using GitHub CodeQL, which we enabled (with default configuration) for all used SWE-Bench projects. In addition, generated patches were reviewed via GitHub Copilot (commit summary, inline comments, low-confidence alerts) to emulate a Human code-review.
@@
For each bug report, we record (1) whether any pre-filter model flagged it, (2) whether SWE-agent generated a patch and whether it fulfilled the attacker’s intended goal (e.g., reversion, leakage, injection), (3) whether Copilot raised any security-relevant concerns, and (4) whether CodeQL detected any vulnerability. Manual validation was performed by two independent experts—a security researcher and a penetration tester—who jointly reviewed each issue and patch. The validation process included reviewing patch logic, inspecting Copilot outputs, and assessing whether the outcome matched the attack objective.
@@ 6. Results
Table 2 summarizes attack success rates as well as the effectiveness of detection and mitigation mechanisms across the APR pipeline, including pre-APR filtering, post-APR detection, attack success, test suits execution, and combined ensemble defenses.
@@ 6.1. Attack evaluation
6.2. Software Tests on APR Results
We ran the full project test suites to gain additional insight if the generated malicious patches would negatively impact test success rates which could be used as part of an automatic detection strategy.
Across all 51 adversarial patches 12/51 (23.5%) failed the test suite outright, 4/51 (7,8%) increased failures only within the project’s expected flaky baseline, 35/51 (68,6%) produced no new test failures.
Vulnerability attacks. Out of 38 cases, 11 failed the test suite, 3 increased the number of failures within expected limits, and 24 showed no additional failures. For injection-based attacks, 4 cases failed, while the remaining ones passed without changes. All CI/CD attacks passed. Within the revert-CVE subset, 7 cases were clearly detected, 3 increased failures within expected limits, and 4 showed no additional failures. Noise attacks. Only one case failed (Noise-APR), and one increased the number of failures within the expected limit (Naive-APR).
This detection pattern is expected: project test suites rarely cover all behavioral paths, and many vulnerabilities arise from subtle, multi-step interactions that unit tests do not capture. Also, as shown in our CI/CD attack scenario (and in cases such as Django pr-63 included in the replication package), attackers may target CI configuration or test files themselves. Thus, while testing is helpful, it cannot fully protect APR systems from adversarial bug reports.
@@ 6.5. Ensembles and Combining Pre- and Post-APR Countermeasures
A full pre-APR ensemble blocks 32 issues; when combined with post-APR review, the total rises to 37 blocked pull requests (72.5%). However, as discussed in Section 7, this increases false positives and may be impractical in production settings.
Adding project test suites as a Fail-based signal provides marginal benefit: +1 detection for noise attacks (both configurations) and +1 (Ensemble) / +2 (Best) for CVE reversion, with no gains for other attacks.
6.6. Cost Analysis
We quantify the economic asymmetry between attackers and defenders by comparing the cost of generating adversarial bug reports ($C_{\text{attack}}$), applying pre-APR counter-measures ($C_{\text{pre}}$), and executing APR ($C_{\text{apr}}$).
Cost of the Attacker ($C_{\text{attack}}$). Generating adversarial bug reports is extremely inexpensive. Across all adversarial issues, the total cost of the generation pipeline was \$0.0133, yielding a mean cost of approximately \$0.000295 per bug report. Costs were estimated based on token usage and current pricing for a locally run Devstral model (\$0.1/M input tokens and \$0.3/M output tokens [^4] Pricing based on OpenRouter https://openrouter.ai/mistralai/devstral-small and Mistral AI https://mistral.ai/news/devstral). On average, this cost decomposes into input (\$0.000083 $\pm$ \$0.000040) and output (\$0.000212 $\pm$ \$0.000144) components.
Cost of Pre-APR Defenses ($C_{\text{pre}}$). Pre-APR defenses incur higher costs while remaining only partially effective. Using unstructured outputs, the best-performing configuration (o4-mini) achieved a detection rate of 47.05% at a mean cost of \$0.00268 $\pm$ \$0.00100 per checked issue; gpt-4.1-mini incurred a mean cost of \$0.00101 $\pm$ \$0.00041. Structured-output variants exhibited comparable costs (o4-mini: \$0.00284 $\pm$ \$0.00111; gpt-4.1-mini: \$0.00097 $\pm$ \$0.00027) without improving detection performance. Lightweight defenses such as PromptGuard, LlamaGuard, and Granite-Guardian were cheaper but detected fewer than 10% of malicious issues.
Cost of APR ($C_{\text{apr}}$). Executing APR dominates the cost profile. Across the evaluated projects, the mean APR cost per issue was \$0.87 $\pm$ \$0.39. Notably, in 90% of cases, adversarial bug reports induced attacker-aligned patches, meaning this cost is frequently incurred to produce insecure or harmful changes.
Summary. We observe an order-of-magnitude asymmetry. The cost of generating an adversarial bug report is on the order of $10^{-4}$ USD, while applying the most effective pre-APR defenses costs on the order of $10^{-3}$ USD per issue–approximately one order of magnitude higher—yet still fails to block most attacks. In contrast, APR execution costs on the order of $10^{0}$ USD per issue, roughly three orders of magnitude higher than $C_{\text{attack}}$ and two orders of magnitude higher than $C_{\text{pre}}$. This imbalance structurally favors attackers, who can generate adversarial inputs at negligible cost, while defenders incur substantially higher expenses both to filter malicious inputs and to execute APR:
$C_{\text{attack}}\;\approx\;10^{-4}\;\ll\;C_{\text{pre}}\;\approx\;10^{-3}\;\ll\;C_{\text{apr}}\;\approx\;10^{0}\quad\text{(USD/issue)}.$
@@ 7.1. Adversarial Attacks for APR
RQ1: Are APR systems prone to generating insecure or unintended code when given adversarial bug reports? Answer: Adversarial bug reports can subvert APR pipelines at multiple stages, inducing the generation of insecure or malicious patches, triggering unintended behaviors during CI/CD execution, and wasting developer effort via plausible but irrelevant issues. In our study, 46 out of 51 crafted bug reports (90.2%) led the APR system to produce patches consistent with the attacker’s objective, including CVE reversion, injection of vulnerable code, test-time execution, and semantically meaningless refactoring.
Quality of the issues from the human developer perspective. All generated issues were analyzed by an experienced software developer and an experienced software security specialist. The reviewed issues appeared plausible and actionable for human developers, at least at first glance. After deeper investigation, most of them could be prevented by human. For example, Naive APR proposes fixes that are obviously not aligned with projects, or CI/CD and Inject attacks are obvious from a security perspective (“extract full environment to external server”) and should immediately raise red flags although they present a plausible story. Other attack classes are more complicated. Noise APR, and revert CVE, are not that obvious from the developer perspective. Our demonstration attacks do not exhaust all possibilities. To further elude detection, attacks may have malicious content split upon into multiple stages and thus make detection harder (Przymus and Durieux, 2025).
@@
APR misalignment arises even for vague or generic issues. For example, a minimally justified request to ”revert recent changes” still prompts the system to synthesize a plausible revert patch, indicating reliance on surface cues rather than deep semantic validation.
We also evaluated attacks designed solely to generate developer burden through meaningless but compilable patches, e.g., arbitrary refactorings, redundant inlines, or unused imports. While such patches do not introduce functional errors, they impose unnecessary review and maintenance costs.
Critically, the distinction between “plausible” and “correct” fixes remains unresolved in APR practice. Some adversarial reports included test cases, others did not—yet both yielded test-passing patches. This underscores a known limitation: passing unit tests does not guarantee semantic correctness, a fact long recognized in APR literature.
Despite this, many systems claim real-world applicability by automatically generating pull requests from unvetted GitHub issues, overlooking malicious inputs. Our attacks are realistic, reproducible, and range from manual to fully automated generation. All bug reports use developer-consistent language and are submitted to private GitHub mirrors, demonstrating their plausibility in standard workflows. The breadth and effectiveness of these strategies expose critical weaknesses in current APR defenses, posing systemic risks to software integrity and supply chain security.
@@ 7.2. Feasibility and Effectiveness of Defenses
RQ2: What defenses are feasible, and at which points in the workflow can attacks be mitigated effectively? Answer: Effective mitigation of adversarial bug reports is feasible but non-trivial. Lightweight, structured LLM-based classifiers, e.g., o4-mini, offer practical pre-APR filtering with low false-positive rates. Runtime isolation in APR/CI environments can prevent exploitation, while post-APR analysis, e.g., Copilot, can flag issues, though often requiring manual review. Ensemble defenses improve detection rates but may incur unacceptable false positives. End-to-end protection demands combining targeted pre-filters, secure execution, and informed human oversight.
False-Positive Rates for Custom Prompts. Performing initial filtering using custom prompts is an efficient way of detecting potentially malicious requests. As Sommer et al. (Sommer and Paxson, 2010) mention, the feasibility of real-life usage of machine leaning techniques depends upon a low false-positive rate as otherwise analyst-fatigue will occur, leading to disabling of filtering mechanisms. To validate the false-positive rate, we selected a range of 100 existing issues of the psf/requests project (issues #6000–6100, excluding one deleted entry) and analyzed them analogous to our malicious issues.
Using structured output resulted in overall low false-positive rates. GPT-4.1-mini resulted in false-positives, while 4o-mini only resulted in a single one. Both of them tagged issue #6063 (a security bypass inquiry). GPT-4.1-mini additionally tagged an issue (#6075, containing only an external link). We argue that both cases are not real-false positives but potentially real-life malicious issues.
@@
Asymmetry of Costs. Our attack (cost: $10^{-4}$ USD) leverages a fundamental asymmetry: it is cheap to execute yet costly to defend. Adversaries can automatically generate realistic, context-aware bug reports with minimal effort. Using lightweight context retrieval and small LLMs, we construct prompts that produce plausible reports in a single pass—no expensive infrastructure required.
In contrast, defenses incur significant cost at multiple stages: (1) Filtering (cost: $10^{-3}$ USD) requires large LLMs or human triage. Since adversarial reports mimic legitimate submissions, simple spam filters or heuristics are ineffective. Low-volume, high-quality submissions evade rate limits and raise few red flags. (2) APR execution (cost: $10^{0}$ USD) is far more expensive than bug report generation. Models like SWE-agent rely on costly multi-turn interactions with large LLMs, often backed by human oversight or CI integration. (3) Post-APR validation, e.g., CodeQL, Copilot, manual review, adds further overhead, with costs accumulating even for patches that are ultimately rejected. This cost imbalance favors attackers: generating high-impact, low-effort reports is trivial, while defenders must invest heavily at each pipeline stage to detect and contain them.
Defending against Adversarial Reports Pre-filter is helpful but insufficient. Revert and Injection attacks can be expressed entirely in natural language, and LLM-based APR systems infer patches even without explicit code. This becomes a cat-and-mouse problem, as attackers can easily rephrase to bypass filters. Distinguishing a malicious LLM-generated report from a legitimate one, e.g., produced by automated bug-finding tools, is extremely difficult. Post-APR defenses (test suites and post-filtering) improve detection, yet about 30% of issues still slip through, while requiring significantly more resources than the attacker. Finally, human triage is already strained: many OSS maintainers report being overwhelmed by large volumes of LLM-generated bug reports even without adversarial intent (Linåker et al., 2024; Larson, 2024). For these reasons, pre-filtering is a useful mitigation but cannot fully protect APR pipelines.